You need 340 mL of a 45% alcohol solution. On hand, you have a 5% alcohol mixture and a 90% alcohol mixture. How much of each mixture will you need to obtain the desired solution?
Answer by mananth(16946)

percent ---------------- quantity
Alcohol type I 5.00% ---------------- x ml
Alcohol type II 90.00% ------ 340 - x ml
Mixture 45.00% ---------------- 340
Total 340 ml
5.00% x + 90.00% ( 340 - x ) = 45.00% * 340
5 x + 90 ( 340 - x ) = 15300
5 x + 30600 - 90 x = 15300
5 x - 90 x = 15300 - -30600
-85 x = -15300
/ -85
x = 180 ml 5.00% Alcohol type I
160 ml 90.00% Alcohol type II

m.ananth@hotmail.ca